OpenStudy (anonymous):

The value of b for which the equation x^2+bx-1=0 and x^2 + x + b=0 have one common root?

OpenStudy (anonymous):

zero works by inspection

OpenStudy (anonymous):

wat do u mean @satellite73

OpenStudy (anonymous):

if you put \(b=0\) you get \[x^2-1\] whose zeros are 1 and -1 and also \(x^2+x\) whose zeros are 0 and -1

OpenStudy (anonymous):

they share one root, namely \(-1\)

OpenStudy (anonymous):

the options are sqrt 2 isqrt3 i sqrt5

OpenStudy (anonymous):

ooooh ok go with \(i\sqrt{3}\)

OpenStudy (anonymous):

hw...???

OpenStudy (anonymous):

solve four equations \[-b\pm\sqrt{b^2+4}=-1\pm\sqrt{1-4b}\] one will give you \(b-0\) which is the answer i like two are impossible one has solution \(\pm i\sqrt{3}\)

OpenStudy (shubhamsrg):

let common root be y, then y ^2 +by - 1=0 = y^2 + y +b =>by -1 = y +b =>b = (y+1)/(y-1) now (y -1)(y+1) = -by (from (1)) and y + 1 = -b/y (from (2)) =>y-1= y^2 y^2 - y +1 = 0 solve for y and then for b..